A sequence of terms is defined as aj = 1, an = (5 - n)an-1. Calculate 2n=1 an.
Answer #1

e 1 = 1 an = 15-n) an- 92 = (5-2) a aula 3a = 3x1 =3 A3 = 15-3) 03-1= 20 = 2x3 = 6 QA (5-4) A. 1x 2 = 1x 6 = 6 95=(5-5) 94 =